OpenStudy (anonymous):

can someone please help??? I don't understand this!!?? :( Barry has been watching the geese that live in his neighborhood. The number of geese changes each week. n f(n) 1 56 2 28 3 14 4 7 Which function best shows the relationship between n and f(n)? f(n) = 28(0.5)^n f(n) = 56(0.5)^n-1 f(n) = 56(0.5)^n f(n) = 112(0.5)^n-1

OpenStudy (tkhunny):

Maybe something that starts at 56 and cuts in half each week?

OpenStudy (phi):

If you are totally lost, try each choice with n=1 according to your table, when n is 1, the answer should be 56 can you figure out 28(0.5)^n when n is 1? in other words, what is \[ 28 \cdot 0.5^1 = 28 \cdot \left(\frac{1}{2}\right)^1 =?\]

OpenStudy (anonymous):

it would be 24

OpenStudy (phi):

ooh, 28/2 = 14 But that is still not 56. so the first choice is not correct.

OpenStudy (anonymous):

oh yeah that is what I mean

OpenStudy (phi):

can you try the 2nd choice?

OpenStudy (anonymous):

with n=1 again????

OpenStudy (phi):

I would write 0.5 as ½ with n=1 then n-1 is 0 so you simplify \[ 56 \cdot \left(\frac{1}{2} \right)^0 \]

OpenStudy (phi):

This one requires you to *remember* anything to the zero power is 1 so 2^0 is 1 3^0 is 1 pi^0 is 1 it always works (except for 0^0 which is not defined)

OpenStudy (anonymous):

because 1/2 times 0 = 0 and 0 times 56 =0 so its not the 2nd choice

OpenStudy (phi):

but we are not doing 2*0 we are doing 2^0 that is a totally different thing. re-read what I posted up above.

OpenStudy (phi):

*remember* anything to the zero power is 1 so 2^0 is 1 3^0 is 1 pi^0 is 1 it always works (except for 0^0 which is not defined)

OpenStudy (anonymous):

oh yeah sorry :) so it would be 28 so not the answer

OpenStudy (phi):

go slowly \[ 56 \cdot \left(\frac{1}{2} \right)^0 \] replace (½)^0 with what ?

OpenStudy (phi):

using *remember* anything to the zero power is 1

OpenStudy (anonymous):

oh sorry I was really confused I replace 1/2^0 with 1 so the answer is 56 which is the answer

OpenStudy (anonymous):

but now I get it if anything has a power of 0 it is automatically 1

OpenStudy (phi):

it is a possible answer. we should check one more value just in case they are trying to trick us. let's try n=2. the table says we should get 28 what is \[ 56 \cdot \left(\frac{1}{2}\right)^{n-1} \] when n is 2 ?

OpenStudy (phi):

*** if anything has a power of 0 it is automatically 1*** with the except of 0^0 which is not defined. we could make sense of it, but not now.

OpenStudy (phi):

replace the n with 2 and figure out 2-1 for the exponent

OpenStudy (anonymous):

1/2^2-1 I would get 1/2^1 which stays 1/2 56*I/2 equals 28

OpenStudy (anonymous):

so its not the answer

OpenStudy (princeharryyy):

2nd is the answer @MilenaSaeger

OpenStudy (anonymous):

I thought its not. im confused now

OpenStudy (princeharryyy):

it's sure @MilenaSaeger

OpenStudy (anonymous):

can u explain why

OpenStudy (princeharryyy):

there's no confusion in it ... ))

OpenStudy (princeharryyy):

hmmm ohk ))

OpenStudy (princeharryyy):

f(1) = 56*(1/2)^(1-1) =56*(1/2)^0 = 56*1 =56; f(2) = 56*(1/2)^(2-1) = 56*(1/2)^1 = 28; f(3) =(56)*(1/2)^(3-1)=56*(1/2)^2 = 56*(1/4) =14 and so on ... hence true ))) :p

OpenStudy (princeharryyy):

@MilenaSaeger :)

OpenStudy (princeharryyy):

????/ r u dere ?????////// @MilenaSaeger

OpenStudy (anonymous):

sorry I just had to walk one of my dogs. oh ok well, thank you two @phi and @princeharryyy

OpenStudy (princeharryyy):

u r welcome :p

OpenStudy (anonymous):

can u help me with one more??

OpenStudy (anonymous):

Zoe keeps track of the miles per gallon her car gets per week. She has accumulated the following data: (1, 24), (2, 24.38), (3, 24.76), (4, 25.14) What is the common difference or ratio? The common difference is 0.38. The common difference is 0.76. The common ratio is 0.99. The common ratio is 1.02. I think it is D

OpenStudy (princeharryyy):

0.38

OpenStudy (princeharryyy):

for increase of 1 in x coordinate y increases by 0.38

OpenStudy (anonymous):

oh ok thank u

OpenStudy (princeharryyy):

or u can say for every x, f(x) that is y increases by 0.38

OpenStudy (princeharryyy):

but how did u got 1.02

OpenStudy (princeharryyy):

@MilenaSaeger

OpenStudy (anonymous):

idk I guessed cause x increased by 1 and y was multiplied by 2. oh nvr mind it only multiplied by 2 once. all the other times it just added 0.38

OpenStudy (anonymous):

i just kinda guessed and used a little common sense

OpenStudy (princeharryyy):

ohk ))))

OpenStudy (princeharryyy):

a bit wrong )))) no big deal

OpenStudy (anonymous):

ok thank u

OpenStudy (princeharryyy):

welcome :)